On October 14, 2012, Fran Co. sold $200,000 of merchandise to Bill Co. Bill gave Fran its 15%, 120-day note for the total.
The amount of accrued interest recognized in an adjusting entry on December 31, 2012 is:
$ 6,500
$ 30,000
$ 6,250
$ 10,000
None of the above
Answer #1

Answer: $6500

Calculation of Interest

Days in October = 31-14 = 17 days

Days in November = 30 days

Days in December = 31 days

Total Days = 17+30+31 = 78 days

Interest calculation = 200000 x 15% x 78/360 = $6500